# Multiscale Training of Convolutional Neural Networks

### v1

@@ Abstract
Convolutional Neural Networks (CNNs) are the backbone of many deep learning methods, but optimizing them remains computationally expensive. To address this, we explore multiscale training frameworks and mathematically identify key challenges, particularly when dealing with noisy inputs. Our analysis reveals that in the presence of noise, the gradient of standard CNNs in multiscale training may fail to converge as the mesh-size approaches to 00, undermining the optimization process. This insight drives the development of Mesh-Free Convolutions (MFCs), which are independent of input scale and avoid the pitfalls of traditional convolution kernels. We demonstrate that MFCs, with their robust gradient behavior, ensure convergence even with noisy inputs, enabling more efficient neural network optimization in multiscale settings. To validate the generality and effectiveness of our multiscale training approach, we show that (i) MFCs can theoretically deliver substantial computational speedups without sacrificing performance in practice, and (ii) standard convolutions benefit from our multiscale training framework in practice.
@@ 1 Introduction
In this work, we consider the task of learning a functional $y({\bf x})=\phi(u({\bf x}))$, where ${\bf x}$ is a position (in 2D ${\bf x}=({\bf x}_{1},{\bf x}_{2})$ and in 3D ${\bf x}=({\bf x}_{1},{\bf x}_{2},{\bf x}_{3})$), $u({\bf x})\in{\cal U}$ and $y({\bf x})\in{\cal Y}$ are families of functions. To this end, we assume to have discrete samples from ${\cal U}$ and ${\cal Y}$, $({\bf u}_{i}^{h}=u_{i}({\bf x}_{h}),{\bf y}^{h}_{i}=\phi(u_{i}({\bf x}_{h}))),\ \ i=1,\ldots,M$ associated with some resolution $h$. A common approach to learning the function is to parameterize the problem, typically by a deep network, and replace $\phi$ with a function $f(\cdot,\cdot)$ that accepts the vector ${\bf u}^{h}$ and learnable parameters ${\bm{\theta}}$ which leads to the problem of estimating ${\bm{\theta}}$ such that
${\bf y}^{h}_{i}\approx f({\bf u}^{h}_{i},{\bm{\theta}}),\quad i=1,\ldots,M.$ (1)
To evaluate ${\bm{\theta}}$, the following stochastic optimization problem is formed and solved:
$\min_{{\bm{\theta}}}\ {\mathbb{E}}_{{\bf u}^{h},{\bf y}^{h}}{\ell}\left(f({\bf u}^{h},{\bm{\theta}}),{\bf y}^{h}\right),$ (2)
Where $\ell(\cdot,\cdot)$ is a loss function, typically mean square error. Standard approaches use variants of stochastic gradient descent (SGD) to estimate the loss and its gradient for different samples of $({\bf u}^{h},{\bf y}^{h})$. In deep learning with convolutional neural networks, the parameter ${\bm{\theta}}$ (the convolutional weights) has identical dimensions, independent of the resolution. Although SGD is widely used, its computational cost can become prohibitively high as the mesh-size $h$ decreases, especially when evaluating the function $f$ on a fine mesh for many samples ${\bf u}_{i}^{h}$. This challenge is worsened if the initial guess ${\bm{\theta}}$ is far from optimal, requiring many costly iterations, for large data sizes $M$. One way to avoid large meshes is to use small crops of the data where large images are avoided, however, this can degrade performance, especially when a large receptive field is required for learning (Araujo et al., 2019)
Background and Related Work. Computational cost reduction can be achieved by leveraging different resolutions, a concept foundational to multigrid and multiscale methods. These methods have a long history of solving partial differential equations and optimization problems (Trottenberg et al., 2001; Briggs et al., 2000; Nash, 2000). Techniques like multigrid (Trottenberg et al., 2001) and algorithms such as MGopt (Nash, 2000; Borzi, 2005) and Multilevel Monte Carlo (Giles, 2015; 2008; Van Barel & Vandewalle, 2019) are widely used for optimization and differential equations.
@@
Our approach. This work falls into the second category of multiscale training. We study multiscale algorithms that use coarse meshes to approximate high-resolution quantities, particularly the gradients of network parameters. Computing gradients on coarse grids is significantly cheaper than on fine grids, as noted in Shi & Cornish (2021). However, for efficient multiscale training, parameters on coarse and fine meshes must have "similar meaning," implying that both the loss and gradient on coarse meshes should approximate those on fine meshes. Specifically, the loss and gradient with respect to parameters should converge to a finite value as $h\rightarrow 0$. In this work, we show that standard CNN gradients may not converge as the mesh size $h$ approaches 00, suggesting CNNs under-utilize multiscale training. This motivates our development of mesh-free convolution kernels, whose values and gradients converge as $h\rightarrow 0$. Our approach builds on Differential-Operator theory (Wong, 2014) to create a family of learnable, mesh-independent convolutions for multiscale learning, resembling Fourier Neural Operators (FNO) (Li et al., 2020a) but with further expressiblity.
@@ 2 Multiscale Stochastic Gradient Descent
We now present the standard approach of training CNNs, identify its major computational bottleneck, and propose a novel solution called Multiscale Stochastic Gradient Descent (Multiscale-SGD).
Standard Training of Neural Networks. Suppose that we use a gradient descent-based method to train a CNN, with input resolution $h$ [^1] In this paper, we define resolution $h$ as the pixel size on a 2D uniform meshgrid, where smaller $h$ indicates higher resolution. For simplicity, we assume the same $h$ across all dimensions, though different resolutions can be assigned per dimension. with trainable parameters ${\bm{\theta}}$. The $k$-th iteration reads:
${\bm{\theta}}_{k+1}={\bm{\theta}}_{k}-\mu_{k}{\mathbb{E}}\left[{\bf g}({\bf u}^{h},{\bf y}^{h},{\bm{\theta}}_{k})\right],$ (3)
where $\ell$ is some loss function (e.g., the mean-squared-error function), and the gradient of the loss with respect to the parameters is ${\bf g}({\bf u}^{h},{\bf y}^{h},{\bm{\theta}})={\bm{\nabla}\,}\ell\left(f({\bf u}^{h},{\bm{\theta}}_{k}),{\bf y}^{h}\right)$. The expectation $\mathbb{E}$ is taken with respect to the input-label pairs $({\bf u}^{h},{\bf y}^{h})$. Evaluating the expected value of the gradient can be highly expensive. To understand why, consider the estimation of the gradient obtained via the average over ${\bf g}$ with a batch of $N$ samples:
@@
This approximation results in an error in the gradient. Under some mild assumptions on the sampling of the gradient value (see Johansen et al. (2010)), the error can be bounded by:
@@
Figure 2: Illustration of the Full Multiscale Training described in Section 4.
@@ 2.1 Efficient Training with Multiscale Stochastic Gradient Descent
To reduce the cost of the computation of the gradient, we use a classical trick proposed in the context of Multilevel Monte Carlo methods (Giles, 2015). To this end, let $h=h_{1}<h_{2}<...<h_{L}$ be a sequence of mesh step sizes, in which the functions $u$ and $y$ are discretized on. We can easily sample (or coarsen) $u$ and $y$ to some mesh $h_{j},1\leq j\leq L$. We consider the following identity, based on the telescopic sum and the linearity of the expectation:
${\mathbb{E}}\left[{\bf g}^{h_{1}}({\bm{\theta}})\right]={\mathbb{E}}\left[{\bf g}^{h_{L}}({\bm{\theta}})\right]\ +\ {\mathbb{E}}\ \left[{\bf g}^{h_{L-1}}({\bm{\theta}})-{\bf g}^{h_{L}}({\bm{\theta}})\right]+\ \ldots\ +\ {\mathbb{E}}\left[{\bf g}^{h_{1}}({\bm{\theta}})-{\bf g}^{h_{2}}({\bm{\theta}})\right],$ (6)
where for shorthand we define the gradient of ${\bm{\theta}}$ with resolution $h_{j}$ by ${\bf g}^{h_{j}}({\bm{\theta}})={\bf g}({\bf u}^{h_{j}},{\bf y}^{h_{j}},{\bm{\theta}})$.
The core idea of our Multiscale Stochastic Gradient Descent (Multiscale-SGD) approach, is that the expected value of each term in the telescopic sum is approximated using a different batch of data with a different batch size. This concept is demonstrated in Figure 1 and it can be summarizes as:
${\mathbb{E}}\left[{\bf g}^{h_{1}}({\bm{\theta}})\right]$$\approx$ ${\frac{1}{N_{L}}}\sum_{i}{\bf g}^{h_{L}}_{i}({\bm{\theta}})\ +\ {\frac{1}{N_{L-1}}}\sum_{i}\left({\bf g}^{h_{L-1}}_{i}({\bm{\theta}})-{\bf g}^{h_{L}}_{i}({\bm{\theta}})\right)+\ \ldots$
$+$ ${\frac{1}{N_{1}}}\sum_{i}\left({\bf g}^{h_{1}}_{i}({\bm{\theta}})-{\bf g}^{h_{2}}_{i}({\bm{\theta}})\right).$
To understand why this concept is beneficial, we analyze the error obtained by sampling each term in Section 2.1. Evaluating the first term in the sum requires evaluating the function ${\bf g}$ on the coarsest mesh (i.e., lowest resolution) using downsampled inputs. Therefore, it can be efficiently computed, while utilizing a large batch size, $N_{L}$. Thus, following Equation 5, the approximation error of the first term in the Section 2.1 can be bounded by:
$\left\|{\mathbb{E}}\left[{\bf g}^{h_{L}}({\bm{\theta}})\right]-{\frac{1}{N_{L}}}\sum_{i}{\bf g}^{h_{L}}_{i}({\bm{\theta}})\right\|^{2}\leq{\frac{C}{\sqrt{N_{L}}}}.$ (8)
Following this step, we need to evaluate the terms of the form
@@
Similarly, this step can be computed by resampling, with a batch size $N_{j}$:
$\hat{{\bf r}}_{j}={\frac{1}{N_{j}}}\sum_{i}\left({\bf g}^{h_{j-1}}_{i}({\bm{\theta}})-{\bf g}^{h_{j}}_{i}({\bm{\theta}})\right),$ (10)
for $j=1,\ldots,L-1$. The key question is: what is the error in approximating ${\bf r}_{j}$ by the finite sample estimate $\hat{{\bf r}}_{j}$? Previously, we focused on error due to sample size. However, note that the exact term ${\bf r}_{j}$ is computed by evaluating ${\bf g}$ on two resolutions of the same samples and subtracting the results.
@@
for some constant $B>0$ and $p>0$, both independent of the pixel-size $h_{j}$. Then, we can bound the error of approximating ${\bf r}_{j}$ by $\hat{{\bf r}}_{j}$, as follows:
@@
Note that, under the assumption that Equation 11 holds, the gradient approximation error between different resolutions decreases as the resolution increases (i.e., $h\rightarrow 0)$. Indeed, sum of the gradient approximation error between subsequent resolutions (where each is defined in Equation 9), where the approximation is obtained from the telescopic sum in Section 2.1, can be bounded by:
@@
Following our Multiscale-SGD approach in Section 2.1, the sample size needed on the finest mesh is reduced by a factor of $4^{L}$ from the original $N_{L}$ while maintaining the same error order, leading to significant computational savings. For instance, Table 3 shows that using 4 mesh resolution levels achieves an order of magnitude in computational savings compared to a single high-resolution gradient approximation.
Beyond these savings, Multiscale-SGD is easy to implement. It simply requires computing the loss at different input scales and batches, which can be done in parallel. Since gradients are linear, the loss gradient naturally yields Multiscale-SGD. The full algorithm is outlined in Algorithm 1.
@@
Set batch size to $N_{L}$ and sample, $N_{L}$ samples of ${\bf u}^{h_{1}}$ and ${\bf y}^{h_{1}}$
Pool ${\bf u}^{h_{L}}={\bf R}^{h_{1}}_{h_{L}}{\bf u}^{h_{1}},\quad{\bf y}^{h_{L}}={\bf R}^{h_{1}}_{h_{L}}{\bf y}^{h_{1}}$
@@
for $j=1,...,L$ (in parallel) do
     Set batch size to $N_{j}$ and sample, $N_{j}$ samples of ${\bf u}^{h_{1}}$ and ${\bf y}^{h_{1}}$
     Pool ${\bf u}^{h_{j}}={\bf R}^{h_{1}}_{h_{j}}{\bf u}^{h_{1}},\quad{\bf y}^{h_{j}}={\bf R}^{h_{1}}_{h_{j}}{\bf y}^{h_{1}}$ and ${\bf u}^{h_{j-1}}={\bf R}^{h_{1}}_{h_{j-1}}{\bf u}^{h_{1}},\quad{\bf y}^{h_{j-1}}={\bf R}^{h_{1}}_{h_{j-1}}{\bf y}^{h_{1}}$
     Compute the losses $\ell({\bf u}^{h_{j}},{\bf y}^{h_{j}},{\bm{\theta}})$ and $\ell({\bf u}^{h_{j-1}},{\bf y}^{h_{j-1}},{\bm{\theta}})$

### v2

# Multiscale Stochastic Gradient Descent: Efficiently Training Convolutional Neural Networks

@@ Abstract
Stochastic Gradient Descent (SGD) is the foundation of modern deep learning optimization but becomes increasingly inefficient when training convolutional neural networks (CNNs) on high-resolution data. This paper introduces Multiscale Stochastic Gradient Descent (Multiscale-SGD), a novel optimization approach that exploits coarse-to-fine training strategies to estimate the gradient at a fraction of the cost, improving the computational efficiency of SGD type methods while preserving model accuracy. We derive theoretical criteria for Multiscale-SGD to be effective, and show that while standard convolutions can be used, they can be suboptimal for noisy data. This leads us to introduce a new class of learnable, scale-independent Mesh-Free Convolutions (MFCs) that ensure consistent gradient behavior across resolutions, making them well-suited for multiscale training. Through extensive empirical validation, we demonstrate that in practice, (i) our Multiscale-SGD approach can be used to train various architectures for a variety of tasks, and (ii) when the noise is not significant, standard convolutions benefit from our multiscale training framework. Our results establish a new paradigm for the efficient training of deep networks, enabling practical scalability in high-resolution and multiscale learning tasks.
Machine Learning, ICML
Niloufar Zakariaei1∗, Shadab Ahamed1∗, Eldad Haber1, Moshe Eliasof2
1
University of British Columbia, Vancouver, Canada
2University of Cambridge, Cambridge, United Kingdom
nilouzk@student.ubc.ca, shadab@phas.ubc.ca, ehaber@eoas.ubc.ca, me532@cam.ac.uk
* Equal contribution
@@ 1 Introduction
In this work, we consider the task of learning a functional $y({\bf x})=\phi(u({\bf x}))$, where ${\bf x}$ is a position (in 2D ${\bf x}=({\bf x}_{1},{\bf x}_{2})$ and in 3D ${\bf x}=({\bf x}_{1},{\bf x}_{2},{\bf x}_{3})$), $u({\bf x})\in{\cal U}$ and $y({\bf x})\in{\cal Y}$ are families of functions. To this end, we assume to have discrete samples from ${\cal U}$ and ${\cal Y}$, $({\bf u}_{i}^{h}=u_{i}({\bf x}_{h}),{\bf y}^{h}_{i}=\phi(u_{i}({\bf x}_{h}))),\ \ i=1,\ldots,M$ associated with some resolution $h$. A common approach to learning the function is to parameterize the problem, typically by a deep network, and replace $\phi$ with a function $f(\cdot,\cdot)$ that accepts the vector ${\bf u}^{h}$ and learnable parameters ${\bm{\theta}}$ which leads to the problem of estimating ${\bm{\theta}}$ such that, ${\bf y}^{h}_{i}\approx f({\bf u}^{h}_{i},{\bm{\theta}})$ for $i=1,\ldots,M$. To evaluate ${\bm{\theta}}$, the following stochastic optimization problem is solved:
$\min_{{\bm{\theta}}}\ {\mathbb{E}}_{{\bf u}^{h},{\bf y}^{h}}{\ell}\left(f({\bf u}^{h},{\bm{\theta}}),{\bf y}^{h}\right).$ (1)
Here, $\ell(\cdot,\cdot)$ is a loss function, typically the mean square error. Standard approaches use variants of stochastic gradient descent (SGD) to estimate the loss and its gradient for different samples of $({\bf u}^{h},{\bf y}^{h})$. In deep learning with convolutional neural networks, the parameter ${\bm{\theta}}$ (the convolutional weights) has identical dimensions, independent of the resolution. Although SGD is widely used, its computational cost can become prohibitively high as the mesh-size $h$ decreases, especially when evaluating the function $f$ on a fine mesh for many samples ${\bf u}_{i}^{h}$. This challenge is worsened if the initial guess ${\bm{\theta}}$ is far from optimal, requiring many costly iterations, for large data sizes $M$. One way to avoid large meshes is to use small crops of the high resolution image, however, this can degrade performance, especially when a large receptive field is required for learning (Araujo et al., 2019).
Background and related work. Computational cost reduction can be achieved by leveraging different resolutions, a concept foundational to multigrid and multiscale methods. These methods have a long history of solving partial differential equations and optimization problems (Trottenberg et al., 2001; Briggs et al., 2000; Nash, 2000). Techniques like multigrid (Trottenberg et al., 2001) and algorithms such as MGopt (Nash, 2000; Borzi, 2005) and Multilevel Monte Carlo (Giles, 2015, 2008; Van Barel & Vandewalle, 2019) are widely used for optimization and differential equations.
@@
Our approach. This work falls into the second category of multiscale training. We study multiscale algorithms that use coarse meshes to approximate high-resolution quantities, particularly the gradients of network parameters. Computing gradients on coarse grids is significantly cheaper than on fine grids, as noted in (Shi & Cornish, 2021). However, for efficient multiscale training, parameters on coarse and fine meshes must have “similar meaning," implying that both the loss and gradient on coarse meshes should approximate those on fine meshes. Specifically, the loss and gradient with respect to parameters should converge to a finite value as $h\rightarrow 0$. In this work, we show that standard CNN gradients may not converge as the mesh size $h$ approaches 00, suggesting CNNs under-utilize multiscale training. This motivates our development of mesh-free convolution kernels, whose values and gradients converge as $h\rightarrow 0$. Our approach builds on Differential-Operator theory (Wong, 2014) to create a family of learnable, mesh-independent convolutions for multiscale learning, resembling Fourier Neural Operators (FNO) (Li et al., 2020a) but with enhanced expressiblity.
@@ 2 Multiscale Stochastic Gradient Descent
We now present the standard approach for training CNNs, identify its major computational bottleneck, and propose a novel solution called Multiscale Stochastic Gradient Descent (Multiscale-SGD).
Standard training of neural networks. Suppose we use a gradient descent-based method to train a CNN, with input resolution [^1] In this paper, we define resolution $h$ as the pixel size on a 2D uniform meshgrid, where a smaller $h$ indicates a higher resolution. For simplicity, we assume the same $h$ across all dimensions, though different resolutions can be assigned per dimension. $h$ with trainable parameters ${\bm{\theta}}$. At $k^{\text{th}}$ iteration, we have
${\bm{\theta}}_{k+1}={\bm{\theta}}_{k}-\mu_{k}{\mathbb{E}}_{{\bf u}^{h},{\bf y}^{h}}\left[{\bf g}({\bf u}^{h},{\bf y}^{h},{\bm{\theta}}_{k})\right],$ (2)
where the gradient of the loss $\ell$ with respect to ${\bm{\theta}}$ is ${\bf g}({\bf u}^{h},{\bf y}^{h},{\bm{\theta}})={\bm{\nabla}\,}\ell\left(f({\bf u}^{h},{\bm{\theta}}_{k}),{\bf y}^{h}\right)$. Evaluating the expected value of the gradient can be highly expensive. To understand why, consider the estimation of the gradient obtained via the average over ${\bf g}$ with a batch of $N$ samples,
@@
This approximation results in an error in the gradient. Under some mild assumptions on the sampling of the gradient value (see Johansen et al. (2010)), the error can be bounded by,
@@
Figure 2: Illustration of the Full Multiscale training using the Full Multiscale Stochastic Gradient Descent (MSGD) algorithm described in Section 4.
@@ 2.1 Efficient training with Multiscale Stochastic Gradient Descent
To reduce the cost of the computation of the gradient, we use a classical trick proposed in the context of Multilevel Monte Carlo methods (Giles, 2015). To this end, let $h=h_{1}<h_{2}<...<h_{L}$ be a sequence of mesh step sizes, in which the functions $u$ and $y$ are discretized with. We can easily sample (or coarsen) $u$ and $y$ to some mesh $h_{j},1\leq j\leq L$. We consider the following identity, based on the telescopic sum and the linearity of the expectation,
${\mathbb{E}}\left[{\bf g}^{h_{1}}({\bm{\theta}})\right]$ $={\mathbb{E}}\left[{\bf g}^{h_{L}}({\bm{\theta}})\right]+{\mathbb{E}}\left[{\bf g}^{h_{L-1}}({\bm{\theta}})-{\bf g}^{h_{L}}({\bm{\theta}})\right]$
$+\ldots\ +{\mathbb{E}}\left[{\bf g}^{h_{1}}({\bm{\theta}})-{\bf g}^{h_{2}}({\bm{\theta}})\right],$ (5)
where for shorthand we define the gradient of ${\bm{\theta}}$ with resolution $h_{j}$ by ${\bf g}^{h_{j}}({\bm{\theta}})={\bf g}({\bf u}^{h_{j}},{\bf y}^{h_{j}},{\bm{\theta}})$. The core idea of our Multiscale Stochastic Gradient Descent (Multiscale-SGD) approach, is that the expected value of each term in the telescopic sum is approximated using a different batch of data with a different batch size. This concept is demonstrated in Figure 1 and it can be summarized as,
${\mathbb{E}}\left[{\bf g}^{h_{1}}({\bm{\theta}})\right]$ $\approx{\frac{1}{N_{L}}}\sum_{i}{\bf g}^{h_{L}}_{i}({\bm{\theta}})$
$+{\frac{1}{N_{L-1}}}\sum_{i}\left({\bf g}^{h_{L-1}}_{i}({\bm{\theta}})-{\bf g}^{h_{L}}_{i}({\bm{\theta}})\right)$
$+\ldots+{\frac{1}{N_{1}}}\sum_{i}\left({\bf g}^{h_{1}}_{i}({\bm{\theta}})-{\bf g}^{h_{2}}_{i}({\bm{\theta}})\right).\vspace{-0.5cm}$ (6)
To understand why this concept is beneficial, we analyze the error obtained by sampling each term in Section 2.1. Evaluating the first term in the sum requires evaluating the function ${\bf g}$ on the coarsest mesh (i.e., lowest resolution) using downsampled inputs. Therefore, it can be efficiently computed, while utilizing a large batch size, $N_{L}$. Thus, following Equation 4, the approximation error of the first term in the Section 2.1 can be bounded by,
$\left\|{\mathbb{E}}\left[{\bf g}^{h_{L}}({\bm{\theta}})\right]-{\frac{1}{N_{L}}}\sum_{i}{\bf g}^{h_{L}}_{i}({\bm{\theta}})\right\|_{2}\leq{\frac{C}{\sqrt{N_{L}}}}.$ (7)
Following this step, we need to evaluate the terms of the form,
@@
Similarly, this step can be computed by resampling, with a batch size $N_{j}$,
$\hat{{\bf r}}_{j}={\frac{1}{N_{j-1}}}\sum_{i}\left({\bf g}^{h_{j-1}}_{i}({\bm{\theta}})-{\bf g}^{h_{j}}_{i}({\bm{\theta}})\right),$ (9)
for $j=2,\ldots,L-1$. The key question is: what is the error in approximating ${\bf r}_{j}$ by the finite sample estimate $\hat{{\bf r}}_{j}$? Previously, we focused on errors due to sample size. However, note that the exact term ${\bf r}_{j}$ is computed by evaluating ${\bf g}$ on two resolutions of the same samples and subtracting the results.
@@
for some constant $B>0$ and $p>0$, both independent of the pixel-size $h_{j}$. Then, we can bound the error of approximating ${\bf r}_{j}$ by $\hat{{\bf r}}_{j}$ as,
@@
Note that, under the assumption that Equation 10 holds, the gradient approximation error between different resolutions decreases as the resolution increases (i.e., $h\rightarrow 0)$. Indeed, the sum of the gradient approximation error between subsequent resolutions (where each is defined in Equation 8), where the approximation is obtained from the telescopic sum in Section 2.1, can be bounded by,
@@
Following our Multiscale-SGD approach in Section 2.1, the sample size needed on the finest mesh is reduced by a factor of $4^{L}$ from the original $N_{L}$ while maintaining the same error order, leading to significant computational savings. For instance, Table 3 shows that using 4 mesh resolution levels achieves orders of magnitude in computational savings compared to a single high-resolution gradient approximation. Beyond these savings, Multiscale-SGD is easy to implement. It simply requires computing the loss at different input scales and batches, which can be done in parallel. Since gradients are linear, the loss gradient naturally yields Multiscale-SGD. The full algorithm is outlined in Algorithm 1.
@@
  Set batch size to $N_{L}$ and sample $N_{L}$ samples of ${\bf u}^{h_{1}}$ and ${\bf y}^{h_{1}}$
  Pool ${\bf u}^{h_{L}}={\bf R}^{h_{1}}_{h_{L}}{\bf u}^{h_{1}}$,  ${\bf y}^{h_{L}}={\bf R}^{h_{1}}_{h_{L}}{\bf y}^{h_{1}}$
@@
  for $j=1,\dots,L$ in parallel do
     Set batch size to $N_{j}$ and sample $N_{j}$ samples of ${\bf u}^{h_{1}}$ and ${\bf y}^{h_{1}}$
     Pool ${\bf u}^{h_{j}}={\bf R}^{h_{1}}_{h_{j}}{\bf u}^{h_{1}}$,  ${\bf y}^{h_{j}}={\bf R}^{h_{1}}_{h_{j}}{\bf y}^{h_{1}}$
     Compute losses:  $\ell({\bf u}^{h_{j}},{\bf y}^{h_{j}},{\bm{\theta}})$ and $\ell({\bf u}^{h_{j-1}},{\bf y}^{h_{j-1}},{\bm{\theta}})$